# KNARsack: Teaching Neural Algorithmic Reasoners to Solve Pseudo-Polynomial Problems

### v1

@@ Appendix B Other pseudo-polynomial problems
Appendix C NAR construction implementation
@@
15 # Edge length encoding construction is omitted for brevity, see Appendix E.
@@
Appendix D NAR reconstruction implementation
@@
31 # Edge length encoding construction is omitted for brevity, see Appendix E.
@@
Listing 2: Python-style pseudocode for NAR reconstruction for Knapsack. The hints discussed in the main text (§2, p. 3) are presented in L7-L14. Out of those, the only less standard is alteration_type. It is similar to the phase hint found in the other algorithms implemented in CLRS-30, with the exeption that alteration_type is always a sequence of alternating zeros and ones. Contrary to construction, where output was obtained from stacked hints, here it is decoded from the final latent embeddings, as standard for CLRS-30.
@@
Appendix E Edge length encoding
@@
Appendix F Homogeneous NAR construction
In Figure 9, we observe that for the regular processor the bottom-right scalar predictions appear lighter than expected, indicating for out-of-distribution the model is predicting lower DP scalar values than the ground-truth ones. A similar phenomenon was reported by Mirjanic et al. [34] in the context of the Bellman-Ford algorithm, where it was suggested to scale the embeddings by a constant factor $c<1$ at every message-passing step. Differently from them, we address this problem by constraining our processor to be homogenous [8], which enforces the model to be invariant with respect to the scale ($f(\alpha x)=\alpha f(x),\alpha>0$) of the item values. Concretely, this entails removing bias terms from the network, as well as eliminating layer normalisations [35] and gating mechanisms [7, p.5]. Additionally, in each step, it is necessary to encode only homogeneous $\mathrm{dp}_{i,:}$ hints, while applying the decode-only hint mode for other hints. As shown in Figure 10, these modifications improve the correlation between true and predicted DP values, and lead to better generalization of the NAR construction model (Table 4), which enables significantly better overall NAR model performance compared to the no-hint baseline.
@@ Appendix F Homogeneous NAR construction
Appendix G Deterministic reconstruction implementation

### v2

@@ Appendix B Other Pseudo-Polynomial Problems
Appendix C Additional Details on the Experimental Setup
Training and Testing Scales
Training on sizes up to 16 with OOD testing on size 64 is the standard for the CLRS-30 benchmark [2] and is therefore followed by the vast majority of NAR works. Since we ultimately aim to include pseudo-polynomial problems in the benchmark, this emerged as the most natural choice. Given that in our reconstruction model the number of graph nodes equals $N+C+1$, training on $(n=32,C=32)$ is already not memory-feasible within the CLRS-30 framework. If we consider node accuracy alone, $4\times$ generalization is gradually becoming obsolete as OOD test, e.g., for the Bellman-Ford algorithm. However, from our baselines (the default models in CLRS-30), we can see that this is not the case for Knapsack. Additionally, in our case we have two parameters that determine the problem size, while all CLRS-30 algorithms have only one parameter. The size of our DP table is $N\times(C+1)$, so with our OOD experiments on parameters $(n=16,C=64)$, $(n=64,C=16)$, and $(n=32,C=32)$, we effectively test $4\times$ generalization as in CLRS-30, as the number of total elements in the table grows fourfold, and with the additional $(n=64,C=64)$ we test up to $16\times$ generalization.
Runtime and Memory Requirements
Runtime and memory requirements are comparable to the ones reported for CLRS-30 benchmark algorithms. Apart from the homogeneity, we do not modify the computational structure of the GNN layers, so we do not incur any slowdown from the GNN architecture. The only slowdown incurred is for reconstruction, where we do twice the amount of steps a standard model would do (see Figure 8). All experiments were conducted on a single NVIDIA A100 GPU with 40 GB of memory. The longest single training run for any model presented in this work was 44 minutes and 56 seconds. All models consumed less than 10 GB of memory, which allowed us to train 4 different models simultaneously. As a result, we were able to conduct all experiments presented in the paper (across all 3 seeds) in less than 10 hours. Inference time is negligible, measuring less than one second per instance across all configurations.
Adapting Tropical Attention
To enable comparison with our model, we made modifications to the Tropical Attention repository [^2] https://github.com/Baran-phys/Tropical-Attention/ (commit 534ac67). In dataloaders.py, the Knapsack dataset generation was modified to use uniform random values in the range $[0,1]$ for item values instead of integer values from value_range (line 264), and the type signature of set_knapsack_01 was updated to accept float values (line 177). A new greedy_decoding method was added to experiment.py to convert the model’s probability predictions into valid Knapsack solutions. The _eval_one_epoch method (experiment.py, line 205) was extended to integrate greedy decoding for exact match calculation. Exact match ratio was added as an evaluation metric. The model was trained on samples where $n=16$ and $C\leq 16$, and tested on the same size combinations as in our experiment.
Appendix D NAR Construction Implementation
@@
15 # Edge length encoding construction is omitted for brevity, see Appendix F.
@@
Appendix E NAR Reconstruction Implementation
@@
31 # Edge length encoding construction is omitted for brevity, see Appendix F.
@@
Listing 2: Python-style pseudocode for NAR reconstruction for Knapsack. The hints discussed in the main text (§2, p. 3) are presented in L7-L14. Out of those, the only less standard is alteration_type. It is similar to the phase hint found in the other algorithms implemented in CLRS-30, with the exception that alteration_type is always a sequence of alternating zeros and ones. Contrary to construction, where output was obtained from stacked hints, here it is decoded from the final latent embeddings, as standard for CLRS-30.
@@
Appendix F Edge Length Encoding
@@
We further elaborate on this empirically important inductive bias by explaining its motivation. As an example, we use the construction model, which has $C+1$ nodes corresponding to capacities $0,1,\ldots,C$. A similar argument can be straightforwardly applied to the reconstruction model. Recall the DP formula $\mathrm{dp}_{i,c}=\max(\mathrm{dp}_{i-1,c},\mathrm{dp}_{i-1,c-w_{i}}+v_{i})$. Observe that at the $i$-th step, which corresponds to item $i$, the following holds for every node $c$: the DP scalar feature associated with node $c$ must either remain unchanged or be updated based on the value of the DP scalar feature associated with node $c^{\prime}=c-w_{i}$, where $c\geq w_{i}$. We see that $c-c^{\prime}=w_{i}$, meaning that for all nodes $c$, the distance to their corresponding candidate node $c^{\prime}$ is equal to the item weight $w_{i}$. NAR processors are based on GNNs where, along every edge $(i,j)$, a message from node $i$ to node $j$, $m_{ij}$, is computed based on the edge embedding and the embeddings of $i$ and $j$, and these messages are then aggregated across all neighboring nodes [2]. As it is currently implemented, the CLRS-30 employs a positional encoding scheme with scalar real values in $[0,1]$. As a result, the model cannot infer if either of the two nodes is a candidate for the other. The introduction of edge length encoding, which assigns categorical features to edges corresponding to their length (i.e., the absolute difference in capacities), means that, to a large extent, computing the message $m_{ij}$ boils down to checking whether the corresponding distance equals $w_{i}$ or not. Note that our use of the absolute difference $|c-c^{\prime}|$ in the definition of edge length encoding reduces the number of required categories, which is possible because candidates $c^{\prime}<c$ are easily identified based on the existing scalar node positional encoding. The cutoff $M$ is used because it is necessary when employing categorical features, and its value is adjusted to the maximum considered item weight. It is important to note that this argumentation is not limited to the Knapsack problem but can be applied to all pseudo-polynomial DP algorithms. Specifically, the same alignment between edge length encoding and the structure of the pseudo-polynomial DP relation holds for other problems: in Subset Sum and Partition problems, the alignment corresponds to differences of possible sums, while in Minimum Coin Exchange, it corresponds to differences of possible target amounts, to name a few.
Appendix G Homogeneous NAR Construction
In Figure 9, we observe that for the regular processor the bottom-right scalar predictions appear lighter than expected, indicating for out-of-distribution the model is predicting lower DP scalar values than the ground-truth ones. A similar phenomenon was reported by Mirjanic et al. [38] in the context of the Bellman-Ford algorithm, where it was suggested to scale the embeddings by a constant factor $c<1$ at every message-passing step. Differently from them, we address this problem by constraining our processor to be homogeneous [8], which enforces the model to be invariant with respect to the scale ($f(\alpha x)=\alpha f(x),\alpha>0$) of the item values. Concretely, this entails removing bias terms from the network, as well as eliminating layer normalisations [39] and gating mechanisms [7, p.5]. Additionally, in each step, it is necessary to encode only homogeneous $\mathrm{dp}_{i,:}$ hints, while applying the decode-only hint mode for other hints. As shown in Figure 10, these modifications improve the correlation between true and predicted DP values, and lead to better generalization of the NAR construction model (Table 5), which enables significantly better overall NAR model performance compared to the no-hint baseline.
@@ Appendix G Homogeneous NAR Construction
Appendix H Deterministic Reconstruction Implementation